Let A, B and C denotes the subset of a set, S and let 𝐶̅denotes the complement of C in set S.



If (A ∩ C) = (B ∩ C) and (A ∩ 𝐶̅) = (B ∩ 𝐶̅), then prove that (A = B).

1
Expert's answer
2021-12-07T13:11:31-0500

Since (A ∩ C) = (B ∩ C) and (A ∩ 𝐶̅) = (B ∩ 𝐶̅) then,

(B ∩ C) - (A ∩ C)= "\empty" and (A ∩ 𝐶̅) - (B ∩ 𝐶̅)= "\empty"

Therefore, (B ∩ C) - (A ∩ C) = (A ∩ 𝐶̅) - (B ∩ 𝐶̅)

This can also be written as,

(A ∩ C) + (A ∩ 𝐶̅)= (B ∩ C)+ (B ∩ 𝐶̅)

This can be represented using union as follows,  

"(A\cap C)\cup(A\cap C')=(B\cap C)\cup(B\cap C')"

Recognize that the left hand side is A since A = "(A\cap C)\cup(A\cap C')" and the right hand side is B since B = "(B\cap C)\cup(B\cap C')."

This shows that, A=B as required.